OpenStudy (anonymous):

What is the value of y in the equation 2(3y + 6 + 3) = 196 − 16?

OpenStudy (rockycatsrule):

welcome to open study

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

but i really need help

OpenStudy (rockycatsrule):

y = 27

OpenStudy (anonymous):

thx

OpenStudy (rockycatsrule):

welcome any more

OpenStudy (anonymous):

I have a few more can you help

OpenStudy (anonymous):

How many solutions does the equation 3x + 6 = −1 − 3 + 4x have?

OpenStudy (rockycatsrule):

i would say many

OpenStudy (anonymous):

but of these Two None One Infinitely many

OpenStudy (rockycatsrule):

oh ok hang on

OpenStudy (rockycatsrule):

Infinitely many

OpenStudy (anonymous):

Wo!

OpenStudy (anonymous):

thx and i have 2 more

OpenStudy (rockycatsrule):

hit me

OpenStudy (anonymous):

What is the solution to the equation 3x + 2(x − 9) = 8x + x − 14?

OpenStudy (anonymous):

Ham,sters are annoying but i <3 cats

OpenStudy (rockycatsrule):

-1

OpenStudy (anonymous):

What is the value of a in the equation a = 2 + 3a + 6?

OpenStudy (rockycatsrule):

and thx @Liz_Beauty45.

OpenStudy (anonymous):

Lol welcome

OpenStudy (anonymous):

cats are pretty gr8 m8 and suprisingly good at math

OpenStudy (rockycatsrule):

-4

OpenStudy (anonymous):

thx bye dont know how to medal but I faned you